OpenStudy (anonymous):

Which expression is larger? Answer 2(5 – 2) • 3 + 3 2[3 + 3(2)] – (1 – 3)

OpenStudy (anonymous):

the second 1 i think

OpenStudy (cwrw238):

work out whats in the parentheses first next priority is multiplication then addition and subtraction have you heard of PEDMAS?

OpenStudy (anonymous):

oh yea i defiantly forgot and didn't use that :p

OpenStudy (anonymous):

2 (3) x 3 + 3 2 x 3= 6 x 3 + 3 18 + 3 = 21 2[3 + 3(2)] – (1 – 3) 2 (3 + 6) - (1 - 3) 2 x 9 - (1 - 3) 2 x 9 - (-2) 18 - - 2 18 + 2 18 + 2 = 20

OpenStudy (anonymous):

In other words, 2(5 – 2) • 3 + 3 is larger.

OpenStudy (anonymous):

oh okay :p